# Symmetry-enriched topological order and quasi-fractonic behavior in $\mathbb{Z}_{N}$ stabilizer codes

### v1

@@ Appendix E Topological index of BB codes with the toric layout
In this section, we explain how to evaluate the fusion rules of the BB code using the algebraic-geometric approach. Explicitly, we consider the $\mathbf{BB}_{p^{k}}(f,g)$ code with the toric layout. By Theorem 2, this requires us to calculate the topological index of the $\mathbf{BB}_{p^{k}}(f,g)$ code through the BKK theorem.
Figure 6: The topological indices of the $\mathbf{BB}_{p}(f,g)$ code with the toric layout for Case 4 under different parameters. Throughout, we assume $\alpha,\beta\geq 0$ and $a\leq 0$. The pink triangles and cyan segments indicate the Newton polygons of $f_{p}$ and $g_{p}$, respectively, and the yellow shaded regions represent the mixed area in each case. The parameter ranges and corresponding topological indices for each case are displayed above the figures. For the boundary cases shown in the second and fourth figures on the top row, the equality of the BKK theorem does not hold, and the topological indices depend explicitly on the coefficients of $f_{p}$ and $g_{p}$.
@@ E.1.1 Origin shift and axis reflection operations
We note that the topological indexes of $\mathbf{BB}_{N}(f,g)$ codes remain invariant under the following two types transformations.
Firstly, the code is invariant under the transformation
@@ E.1.2 BKK Theorem
We now introduce the Bernstein-Khovanskii-Kushnirenko (BKK) theorem, our main algebraic-geometric tool for evaluating the topological index. We focus on the bivariate polynomial case. To prepare for the theorem, we first define Newton polygons and the concept of mixed area. (The BKK method was first applied to qubit codes in [50] and is generalized here to arbitrary qudits.)
We use the shorthand $X^{\mathbf{j}}$ for $x^{j_{1}}y^{j_{2}}$ with $\mathbf{j}=(j_{1},j_{2})$. The Newton polygon of a bivariate Laurent polynomial
$h=\sum_{\mathbf{j}=(j_{1},j_{2})\in\mathbb{Z}^{2}}c_{\mathbf{j}}X^{\mathbf{j}}$ (105)
@@
If $\operatorname{MV}(\mathscr{P}_{1},\mathscr{P}_{2})>0$, then the equality holds if and only if for all faces $S=S_{1}+S_{2}$ of $\mathscr{P}$, their restricted polynomials $h_{1,S_{1}}$ and $h_{2,S_{2}}$ share no common roots in $(\overline{\mathbb{K}}^{*})^{2}$.
@@ E.2 Topological index
$f=c^{(0)}+c^{(1)}x+c^{(2)}x^{\bar{\alpha}}y^{b},\quad g=d^{(0)}+d^{(1)}x+d^{(2)}x^{a}y^{\bar{\beta}}$ (109)
@@
$f_{p}=c^{(0)}_{p}+c^{(2)}_{p}x^{\bar{\alpha}}y^{b},\quad g_{p}=d^{(0)}_{p}+d^{(2)}_{p}x^{a}y^{\bar{\beta}},$ (111)
@@
$\alpha\beta-ab=0,\quad\left(-\frac{c^{(0)}_{p}}{c^{(2)}_{p}}\right)^{s}=\left(-\frac{d^{(0)}_{p}}{d^{(2)}_{p}}\right)^{t},$ (112)
is satisfied. Here, $s=|a/\gcd(\alpha,a)|$ and $t=|\alpha/\gcd(\alpha,a)|$ arise from the parameterization $\alpha=-tq_{1}$, $\beta=-sq_{2}$, $a=sq_{1}$, and $b=tq_{2}$, with $q_{1},q_{2}\in\mathbb{Z}$. In all other scenarios, the code is topological. Specifically, if $\alpha\beta-ab=0$ but the equality is not met, the code is in a trivial phase with $Q=1$.
If the model is topological, the Newton polygons of $f$ and $g$ are two segments defined by vectors $(-\alpha,b)$ and $(a,-\beta)$. The mixed area is given by the area of the parallelogram spanned by these two vectors. According to the BKK Theorem, the topological index is
@@
$b=\beta=0,\quad\gcd(c^{(0)}_{p}+c^{(1)}_{p}x+c^{(2)}_{p}x^{\alpha},d^{(0)}_{p}+d^{(1)}_{p}x^{a})\neq 1.$ (115)
@@
$f_{p}=c^{(0)}_{p}+c^{(1)}_{p}x+c^{(2)}_{p}x^{\bar{\alpha}}y^{b},\quad g_{p}=d^{(0)}_{p}+d^{(1)}_{p}x+d^{(2)}_{p}x^{a}y^{\bar{\beta}},$ (116)
@@
Appendix F Sample code in M2
In this section, we provide a brief demonstration of basic M2 command lines that are useful for computing the topological order and quasi-fractonic behavior in BB codes. We illustrate this with an example consistent with the main text. The code that produces Gröbner basis is as follows:
@@ Appendix F Sample code in M2
i2: f = 1 + x + 6t*y^5; g = 1 + y + 4x^3*u; -- Define polynomials.
i4: I = ideal(12, tx - 1, yu - 1, f, g); --Define the ideal.
@@
The command for generating a Gröbner basis applies not only to rings, but also to modules, allowing us to directly check the topological condition as follows:
@@
i6: I’= ideal (4, x*t-1, y*u-1)
i7: Q’=S/I -- Define quoitent ring.
i8: epslion = matrix{{f, g}}; partial = g $|$ -f; --Define stabilizer and maps.
i10: gens gb == gens gb partial; -- Check exactness.
o10 = true
The periodicity can then be tested directly in the quotient ring $\mathcal{Q}$. For example, the periodicity along the $x$-direction can be obtained with
@@
i11: Q = Q’/I;
i12: i = 1; g = x; while i > 0 do (if g - 1 == 0 then break i; g = x*g, i = i + 1); i; -- Check the ideal memberships.
o13 = 2
This gives that the periodicity along the $x$-direction is 2.

### v2

@@ Appendix E Topological index of BB codes with the toric layout
In this appendix, we explain how to evaluate the fusion rules of the BB code using the algebraic-geometric approach. Explicitly, we consider the $\mathbf{BB}_{p^{k}}(f,g)$ code with the toric layout. By Theorem 2, this requires us to calculate the topological index of the $\mathbf{BB}_{p^{k}}(f,g)$ code through the BKK theorem.
Figure 6: The topological indices of the $\mathbf{BB}_{p}(f,g)$ code with the toric layout for case 4 under different parameters. Throughout, we assume $\alpha,\beta\geq 0$ and $a\leq 0$. The pink triangles and cyan segments indicate the Newton polygons of $f_{p}$ and $g_{p}$, respectively, and the yellow shaded regions represent the mixed area in each case. The parameter ranges and corresponding topological indices for each case are displayed above the figures. For the boundary cases shown in the second figure on the top row and the first figure on the bottom row, the equality of the BKK theorem does not hold, and the topological indices depend explicitly on the coefficients of $f_{p}$ and $g_{p}$.
@@ E.1.1 Origin shift and axis reflection operations
We note that the topological indexes of $\mathbf{BB}_{N}(f,g)$ codes remain invariant under the following two type transformations.
First, the code is invariant under the transformation
@@ E.1.2 BKK theorem
We now introduce the BKK theorem, our main algebraic-geometric tool for evaluating the topological index. We focus on the bivariate polynomial case. To prepare for the theorem, we first define Newton polygons and the concept of mixed area. (The BKK method was applied to qubit codes in [50] and is generalized here to arbitrary qudits.)
We use the shorthand $\mathcal{X}^{\mathbf{j}}$ for $x^{j_{1}}y^{j_{2}}$ with $\mathbf{j}=(j_{1},j_{2})$. The Newton polygon of a bivariate Laurent polynomial
$h=\sum_{\mathbf{j}=(j_{1},j_{2})\in\mathbb{Z}^{2}}c_{\mathbf{j}}\mathcal{X}^{\mathbf{j}}$ (108)
@@
If $\operatorname{MV}(\mathscr{P}_{1},\mathscr{P}_{2})>0$, then the equality holds if and only if for all faces $S=S_{1}+S_{2}$ of $\mathscr{P}$, their restricted polynomials $h_{1,S_{1}}$ and $h_{2,S_{2}}$ share no common roots in $(\overline{\mathbb{K}}^{\times})^{2}$.
@@ E.2 Topological index
$f=c^{(0)}+c^{(1)}x+c^{(2)}x^{\bar{\alpha}}y^{b},\quad g=d^{(0)}+d^{(1)}y+d^{(2)}x^{a}y^{\bar{\beta}}$ (112)
@@
$f_{p}=c^{(0)}_{p}+c^{(2)}_{p}x^{\bar{\alpha}}y^{b},\quad g_{p}=d^{(0)}_{p}+d^{(2)}_{p}x^{a}y^{\bar{\beta}}.$ (114)
@@
$\alpha\beta-ab=0,\quad\left(-\frac{c^{(0)}_{p}}{c^{(2)}_{p}}\right)^{s}=\left(-\frac{d^{(0)}_{p}}{d^{(2)}_{p}}\right)^{t}.$ (115)
Here, $s=-a/\gcd(\alpha,a)$ and $t=\alpha/\gcd(\alpha,a)$ arise from the parameterization $\alpha=-tq_{1}$, $\beta=-sq_{2}$, $a=sq_{1}$, and $b=tq_{2}$, with $q_{1},q_{2}\in\mathbb{Z}$ (If $\gcd(\alpha,a)=0$, then $s$ and $t$ are defined by $\beta$ and $b$). In all other scenarios, the code is topological. Specifically, if $\alpha\beta-ab=0$ but the equality is not met, the code is in a trivial phase with $Q=0$.
If the model is topological, the Newton polygons of $f_{p}$ and $g_{p}$ are two segments defined by vectors $(-\alpha,b)$ and $(a,-\beta)$. The mixed area is given by the area of the parallelogram spanned by these two vectors. According to the BKK Theorem, the topological index is
@@
$b=\beta=0,\quad\gcd(c^{(0)}_{p}+c^{(1)}_{p}x+c^{(2)}_{p}x^{\bar{\alpha}},d^{(0)}_{p}+d^{(2)}_{p}x^{a})\neq 1.$ (118)
@@
$f_{p}=c^{(0)}_{p}+c^{(1)}_{p}x+c^{(2)}_{p}x^{\bar{\alpha}}y^{b},\quad g_{p}=d^{(0)}_{p}+d^{(1)}_{p}y+d^{(2)}_{p}x^{a}y^{\bar{\beta}},$ (119)
@@
Appendix F Sample codes in M2
In this appendix, we provide a brief demonstration of basic M2 command lines that are useful for computing the topological order and quasifractonic behavior in BB codes. We illustrate this with an example consistent with the main text. The code that produces Gröbner basis is as follows:
@@ Appendix F Sample codes in M2
i2: f = 1 + x + 6*t*y^5; g = 1 + y + 4*x^3*u; -- Define polynomials.
i4: I = ideal(12, t*x - 1, y*u - 1, f, g); --Define the ideal.
@@
The command for generating a Gröbner basis applies not only to rings, but also to modules, allowing us to directly check the topological condition as follows (For computational efficiency, we use the graded reverse lexicographic order here):
@@
i1: S = ZZ[u, t, y, x]; --Define the polynomial ring.
i2: I’= ideal (12, x*t-1, y*u-1);
i3: Q’= S/I -- Define quotient ring.
i4: f = 1 + x + 6*t*y^5; g = 1 + y + 4*x^3*u;
i6: eps = matrix{{f, g}}; partial = g $|$ -f; --Define stabilizer and excitation maps.
i8: try (gens gb kernel eps - gens gb partial == 0) else false -- Check exactness.
o8 = true
The periodicity can then be tested directly in the quotient ring $\mathcal{Q}$. For example, the periodicity along the $\hat{x}$ direction can be obtained with
@@
i1: S = ZZ[u, t, y, x];
i2: f = 1 + x + 6*t*y^5; g = 1 + y + 4*x^3*u;
i4: I = ideal(12, t*x - 1, y*u - 1, f, g);
i5: Q = S/I;
i6: i = 1; g = x; while i > 0 do (if g - 1 == 0 then break i; g = x*g, i = i + 1); i -- Check the ideal memberships.
o9 = 2
This gives that the periodicity along the $\hat{x}$ direction is 2.